# Normalized Conditional Mutual Information Surrogate Loss for Deep Neural Classifiers

### v2

@@ V Experiments
To illustrate the effectiveness of NCMI and compare it with some state-of-the-art alternatives, a series of experiments was conducted. Specifically, we conduct experiments on two widely used natural image datasets, namely CIFAR-100 [krizhevsky2009learning] and ImageNet [5206848], as well as two whole-slide image datasets, namely CAMELYON-17 [8447230] and BRACS [brancati2022bracs]. In the tables, NCMI-LP and NCMI-CC denote NCMI evaluated with linear probing (LP) and with centroids comparison (CC), respectively.
@@ V-A Experiments on CIFAR-100
To illustrate the effectiveness of NCMI, we have conducted experiments on models of varying sizes. Specifically, we have selected ResNet-18, ResNet-34, ResNet-50 and ResNet-101 for evaluation, and compare NCMI with respect to 8 benchmark methods namely, CE, LS [szegedy2016rethinking], AntiClass [kat2025anticlasses], Squentropy [hui2023cut], SquareLoss [hui2021evaluation], PolyLoss [leng2022polyloss], SupCon [khosla2020supervised] and Focal Loss [lin2017focal].
For all surrogate losses, we use an SGD optimizer with a momentum of 0.9, a learning rate of 0.1, and a weight decay of 0.0005, along with a batch size of 64. We train the model for 240 epochs, and at epochs 60, 120, and 160, we reduce the current learning rate by a factor of 10. Since SupCon relies on a large batch size to work, we report the results they reported in the paper and the reproduced results under the same setting.
The results are reported in Table˜I. As seen, the models trained by NCMI outperform those trained by the benchmark methods. Importantly, the improvement is consistent across various model sizes.
@@ V-B Experiments on ImageNet
TABLE II: Top-1 and Top-5 validation accuracy on ImageNet for models trained with NCMI and baseline methods. $C$ denotes greedy prediction via centroid comparison; LP denotes linear probing. The best and second-best results are shown in bold and underlined, respectively.
@@ V-C Experiments on Whole Slide Image Dataset
Figure 4: The evolution curves of ResNet-18 on CIFAR-100 under all combinations of NSF and feature centering (enabled/disabled), w.o.C/w.C denote without/with centering, and SM/NSF denote applying softmax/NSF. Shown are the epoch-wise trajectories of (a) CMI, (b) $\Gamma$, (c) NCMI, and (d) accuracy.
@@ V-D Training Cost & Training Stability
NCMI also converges reliably with small batches. Figure˜3 shows how batch size affects the validation accuracy of CE, SupCon, and NCMI. We train ResNet-50 on CIFAR-100 with batch sizes {16, 32, 64, 128, 256, 1024}. Across all batch sizes, NCMI exhibits robust convergence and consistently surpasses CE. Furthermore, we observe that SupCon relies heavily on large batches; reducing the batch size results in a significant decline in accuracy.
@@ VI Ablation Study
To understand how these components affect learning, we visualize the ResNet-18 training trajectories of CMI, $\Gamma$, NCMI, and accuracy on the CIFAR-100 test set in Figure˜4. As shown, the NSF plays a pivotal role in NCMI training: when the softmax function is used to map features to probability distributions, the model fails to converge. Feature centering further stabilizes training. Specifically, with the NSF enabled but without centering, the model collapses around epoch 150; after centering is enabled, it converges properly. Due to the space limit, we further evaluate the effectiveness of NSF and the centering operation by visualizing class clusters and their centers in the appendix [^2] Extended version (with appendix): https://arxiv.org/pdf/2601.02543..
@@ VII Conclusion
Appendix A Extended Ablations: Feature Centering and Normalized Sigmoid (NSF)

### v3

@@ V Experiments
To illustrate the effectiveness of NCMI and compare it with state-of-the-art CE alternatives, a series of experiments was conducted. We conduct experiments on two widely used natural image datasets, namely CIFAR-100 [krizhevsky2009learning] and ImageNet [5206848], as well as two whole-slide image datasets, namely CAMELYON-17 [8447230] and BRACS [brancati2022bracs]. In the tables, NCMI-LP and NCMI-CC denote NCMI evaluated with linear probing (LP) and with centroids comparison (CC), respectively.
@@ V-A Experiments on CIFAR-100
To illustrate the effectiveness of NCMI, we evaluate NCMI on ResNet-18/34/50/101, whose penultimate feature dimensions are 512, 512, 2048, and 2048, respectively, and compare NCMI with respect to 8 benchmark methods namely, CE, LS [szegedy2016rethinking], AntiClass [kat2025anticlasses], Squentropy [hui2023cut], SquareLoss [hui2021evaluation], PolyLoss [leng2022polyloss], SupCon [khosla2020supervised] and Focal Loss [lin2017focal].
For all surrogate losses, we use an SGD optimizer with a momentum of 0.9, a learning rate of 0.1, and a weight decay of 0.0005, along with a batch size of 64. We train the model for 200 epochs, and at epochs 60, 120, and 160, we reduce the current learning rate by a factor of 10. Since SupCon relies on a large batch size to work, we report the results they reported in the paper and the reproduced results under the same setting.
The results are reported in Table˜I. As seen, NCMI yields consistent improvements across all backbones under both evaluation protocols. With linear probing, NCMI-LP achieves the best accuracy for every architecture, improving over CE by $1.50\%$ (ResNet-18), $1.12\%$ (ResNet-34), $0.80\%$ (ResNet-50), and $0.84\%$ (ResNet-101). Moreover, the strongest gains occur for the smaller simplex dimension.
@@ V-B Experiments on ImageNet
TABLE II: Top-1 and Top-5 validation accuracy on ImageNet for models trained with NCMI and baseline methods. $C$ denotes greedy prediction via nearest-centroid comparison; LP denotes linear probing. The best and second-best results are shown in bold and underlined, respectively.
@@
The results are summarized in Table˜II. With ResNet-50, NCMI improves top-1 accuracy by 2.77% over CE.
@@ V-D Training Cost & Training Stability
NCMI also converges reliably with small batch size. Figure˜3 shows how batch size affects the validation accuracy of CE, SupCon, and NCMI. We train ResNet-50 on CIFAR-100 with batch sizes {16, 32, 64, 128, 256, 1024}. Across all batch sizes, NCMI exhibits robust convergence and consistently surpasses CE. Furthermore, we observe that SupCon relies heavily on large batches; reducing the batch size results in a significant decline in accuracy.
@@ VI Ablation Study
To better understand how these components influence training, we further assess the impact of NSF and the centering operation by visualizing the resulting class clusters and their corresponding centers. Due to space constraints, these results are provided in the appendix [^2] Extended version (with appendix): https://arxiv.org/pdf/2601.02543..
@@ VII Conclusion
Appendix A Proof of Theorem˜1
We present the proof of Theorem˜1 in this section.
$\frac{\frac{1}{|\mathcal{D}|}\sum_{y}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}D(\boldsymbol{p_{x}}\|\boldsymbol{s}^{y})}{\frac{1}{|\mathcal{D}|^{2}}\sum_{\boldsymbol{z}\in\mathcal{D}}\sum_{\boldsymbol{x}\in\mathcal{D}}\mathds{1}_{\{\boldsymbol{c_{x}}\neq\boldsymbol{c_{z}}\}}D(\boldsymbol{s^{\boldsymbol{c_{x}}}}\|\boldsymbol{p_{z}})}$
$=$ $\frac{\frac{1}{|\mathcal{D}|}\sum_{y}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}\Big[\sum_{i=1}^{n}\boldsymbol{p_{x}}[i]\ln\frac{\boldsymbol{p_{x}}[i]}{\boldsymbol{s}^{y}[i]}\Big]}{\frac{1}{|\mathcal{D}|^{2}}\sum_{\boldsymbol{z}\in\mathcal{D}}\sum_{\boldsymbol{x}\in\mathcal{D}}\mathds{1}_{\{\boldsymbol{c_{x}}\neq\boldsymbol{c_{z}}\}}\Big[\sum_{i=1}^{n}\boldsymbol{s}^{\boldsymbol{c_{x}}}[i]\ln\frac{\boldsymbol{s}^{\boldsymbol{c_{x}}}[i]}{\boldsymbol{p_{z}}[i]}\Big]}$ (11)
$=$ $\min_{\boldsymbol{q}^{v},v\in[C]}\frac{\frac{1}{|\mathcal{D}|}\sum_{y}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}\Big[\sum_{i=1}^{n}\boldsymbol{p_{x}}[i]\ln\frac{\boldsymbol{p_{x}}[i]}{\boldsymbol{q}^{y}[i]}\Big]}{\frac{1}{|\mathcal{D}|^{2}}\sum_{\boldsymbol{z}\in\mathcal{D}}\sum_{\boldsymbol{x}\in\mathcal{D}}\mathds{1}_{\{\boldsymbol{c_{x}}\neq\boldsymbol{c_{z}}\}}\Big[\sum_{i=1}^{n}\boldsymbol{s}^{\boldsymbol{c_{x}}}[i]\ln\frac{\boldsymbol{q}^{\boldsymbol{c_{x}}}[i]}{\boldsymbol{p_{z}}[i]}\Big]}$ (12)
$=$ $\min_{\boldsymbol{q}^{v},v\in[C]}\frac{\frac{1}{|\mathcal{D}|}\sum_{y}\sum_{\boldsymbol{x}\in\mathcal{D}^{y}}D(\boldsymbol{p_{x}}\|\boldsymbol{q}^{y})}{\frac{1}{|\mathcal{D}|^{2}}\sum_{\boldsymbol{z}\in\mathcal{D}}\sum_{\boldsymbol{x}\in\mathcal{D}}\mathds{1}_{\{\boldsymbol{c_{x}}\neq\boldsymbol{c_{z}}\}}\Big[H(\boldsymbol{p_{x}},\boldsymbol{p_{z}})-H(\boldsymbol{p_{x}},\boldsymbol{q^{c_{z}}})\Big]}$ (13)
Appendix B Extended Ablations: Feature Centering and Normalized Sigmoid (NSF)